Find the distance of two charges with q1 = 94 nC and q2 = 45 μC, the force of the 2 charges is 0.81 N


1
Expert's answer
2022-05-22T14:39:33-0400
"F=k\frac{q_1q_2}{r^2}"

"r=\sqrt{\frac{kq_1q_2}{F}}\\
=\sqrt{\frac{9*10^9*94*10^{-9}*45*10^{-6}}{0.81}}=0.22\:\rm m=22\: cm"